# Evaluating Sample Utility for Data Selection by Mimicking Model Weights

### v2

@@ Appendix A Grad-Mimic Algorithm
We summarize the algorithm steps in Grad-Mimic for computing mimic scores and updating model weights. Our algorithm supports various training paradigms, including both supervised learning and self-supervised learning. We validate its effectiveness through controlled simulation experiments (Sec. 5.1) and large-scale training of CLIP models on web-crawled datasets (Sec. 5.2). Furthermore, our experiments encompass a range of training configurations, such as full-parameter fine-tuning, linear probing, and training models from scratch.
Algorithm 1 Grad-Mimic (Stage 1)
@@
reference model $\theta_{\text{ref}}$, new model $\theta$, dataset $D$, number of training steps $T$, batch size $b$, learning rate $\eta$, temperature $\tau$
normalized mimic scores $\overline{m}_{i,t}$, a well-trained model $\theta_{T}$
Initialize new weights $\theta_{0}$
for $t\in[T]$ do
     uniformly sample a batch $\{s_{i}\}_{i=1}^{b}$ from $D$
     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as a guide
     for $i\in[b]$ do
         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradient
         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|}$triangleright compute mimic score
     end for
     $\overline{m}_{:,t}=\frac{e^{m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{m_{j,t}/\tau}}$triangleright normalize mimic scores
@@
     $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}$triangleright reweight gradients
end for
Appendix B Experimental Details
We provide more details about our training setups and computational resources.
B.1 Simulation Experiments
@@
In the experiments of simulating mislabeled samples (Sec. 5.1), we fine-tune ViT-B/16 models pretrained on the ImageNet-21k dataset [51]. Each dataset is trained with a batch size of 32, a learning rate of 1e-4, the AdamW optimizer, and for 5 epochs. We evaluate Grad-Mimic using various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report the performance at $\tau=0.5$ in the main paper. Results for other temperature values are presented in Table 7. To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets but with different random seeds for weight initialization. These experiments were performed on a Nvidia Tesla A100.
B.2 DataComp Experiments
@@
We adopt the training setup from DataComp [49, 17] and train CLIP models from scratch using contrastive objective over image-caption pairs. Each CLIP model is trained for 5 epochs with a batch size of 4096. The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset. These experiments run on 8 Nvidia Tesla A100s.
@@
We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders. Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [17] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5). We report results at $\tau=0.5$ and $\tau=0.05$ in the main body in Table 3, but full results are presented in Table 6.
@@
Appendix C Ablation Studies
@@
We test Grad-Mimic using different temperature values and compare them to baseline methods (Mini-batch SGD and Grad-Match [24]). The results in the simulation experiments are presented in Table 7. In these datasets, we set the noise level to 0.3 and fine-tune ViT-B/16 model under linear probing configuration. Grad-Mimic outperforms the baseline methods in all temperature settings. Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.
DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
Table 8: Grad-Mimic outperforms other reference model-based methods: Using the last-layer weights of reference model is more effective to prioritize samples for learning while reducing memory usage and computational overhead.
Comparison to Reference Model-Based Approaches.
We compare Grad-Mimic to Rho-Loss [15], a method that prioritizes samples for learning based on excess loss, computed as the difference between the loss on the current model $\ell(\theta_{t};s_{i})$ and the loss on the pretrained reference model $\ell(\theta_{\text{ref}};s_{i})$. We follow our simulation experiment setup and evaluate the performance of fine-tuned models on the testing datasets.
As shown in Table 8, Grad-Mimic consistently outperforms Rho-Loss across all datasets and noise levels. Furthermore, Grad-Mimic offers efficiency advantages. Unlike Rho-Loss, which requires loading the entire reference model and performing additional inference passes, Grad-Mimic computes mimic scores directly from the reference model’s weights. In our experiments, we demonstrate that even using only a subset of the weights (e.g., the last-layer weights) is sufficient, further reducing memory usage and improving computational efficiency.
Figure 7: Grad-Mimic can outperform influence function-based methods: Using model weights as a guide offer advantages on their effectiveness and greater accessibility.
Comparison to Influence Function-based Approaches.
Data influence functions provide a principled way to evaluating the impact of individual training samples on model parameters [52, 53], but their computational cost often limits practical use. Several studies have explored efficient approximations that estimate sample influence by measuring gradient alignment with a validation dataset [54, 55]. We adapt these methods within Grad-Mimic framework by replacing the vector induced by reference model’s weights ($v_{t}$) with gradients computed from validation dataset. We evaluate these adaptations on six image datasets by adding 50% label noise and assess the impact of validation set size by varying the number of samples drawn from each class (1, 5, 10, 15, and 20).
Results in Figure 7 show that while larger validation sets improve data efficiency by enabling model to focus on the influential samples, Grad-Mimic—using pretrained reference weights and their induced direction—consistently outperforms validation dataset-based influence approximations. Moreover, relying on a labeled validation set risk introducing an unwanted dependency. For example, in the out-of-domain Flowers102 dataset, achieving satisfactory performance required over 2000 correctly labeled validation samples—yet still fell short of Grad-Mimic’s results.
@@
Impact of Reference Model Quality.
Finally, we investigate how the quality of the reference model impacts the effectiveness of Grad-Mimic, particularly in scenarios where only a weaker or distilled model is available. To simulate such scenarios, we inject zero-mean Gaussian noise $(0,\sigma^{2})$ into the reference model’s weights. By varying the variance of the added noise, we analyze the performance of models trained with Grad-Mimic against standard training methods that do not leverage a reference model.
We present this analysis in Figure 8. As expected, increasing noise variance degrades the reference model’s performance, which in turn lowers the effectiveness of Grad-Mimic. However, we observe that even under significant noise, models trained with Grad-Mimic still outperform Mini-batch SGD and can even surpass the performance of the noisy reference model itself. This highlights that while a high-quality reference model is ideal, even a weaker proxy can provide valuable guidance and improve training outcomes over traditional baselines.

### v3

@@ Appendix B Grad-Mimic Algorithms
Here, we summarize algorithm steps. Grad-Mimic proceeds in two stages. Stage 1 (Sec. 4.1) computes mimic scores on the fly and updates model weights through re-weighting. Stage 2 (Sec. 4.2) binarizes mimic scores and aggregates utility assessments to obtain the final estimate. Our algorithm supports various training paradigms, including supervised learning and self-supervised learning, as demonstrated in Sec. 5.1 and Sec. 5.2. Our experiments encompass a range of training configurations, including full-parameter fine-tuning, linear probing, and training models from scratch.
Algorithm 1 Online Batch Reweighting (Stage 1)
@@
1:Reference model $\theta_{\text{ref}}$, model being trained $\theta$, dataset $D$, number of training steps $T$, batch size $b$, learning rate $\eta$, temperature parameter $\tau$
2:Normalized mimic scores $\overline{m}_{i,t}$ for each sample $i$ at step $t$
3:Initialize model weights $\theta_{0}$
4:for $t=1$ to $T$ do
5:     Sample a batch $\{s_{i}\}_{i=1}^{b}$ uniformly from $D$
6:     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright Weight difference vector
7:     for $i=1$ to $b$ do
8:         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright Per-sample gradient
9:         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|}$triangleright Mimic score
10:     end for
11:     $\overline{m}_{:,t}\leftarrow\frac{\exp(m_{:,t}/\tau)}{\sum_{j=1}^{b}\exp(m_{j,t}/\tau)}$triangleright Normalize mimic scores
@@
12:     $\theta_{t+1}\leftarrow\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}$triangleright Update weights with reweighted gradients
13:end for
Algorithm 2 Offline Sample Selection (Stage 2)
[…]
2:Aggregated retain probabilities $p_{i}$ for each sample $i$, retained subset $\mathcal{S}$
3:Initialize binary decision matrix $B\leftarrow\mathbf{0}^{n\times T}$
4:for $t=1$ to $T$ do
5:     $m\leftarrow\overline{M}_{:,t}$triangleright Normalized mimic scores at step $t$
[…]
7:         $\tau\leftarrow 1/b$triangleright Threshold based on batch size
8:         $B_{:,t}\leftarrow\mathbf{1}[m>\tau]$
[…]
10:         Fit a two-cluster $k$-means model on $m$triangleright 1D clustering
11:         $B_{:,t}\leftarrow$ cluster assignments
[…]
13:         $\tau\leftarrow k$-th percentile of $m$
14:         $B_{:,t}\leftarrow\mathbf{1}[m\geq\tau]$
15:     end if
16:end for
17:Train a Snorkel LabelModel $\mathcal{LM}$ on $B$
18:$P\leftarrow\mathcal{LM}.\text{predict\_proba}(B)$triangleright Retain/discard probabilities, $P\in\mathbb{R}^{N\times 2}$
19:$p_{i}\leftarrow P_{i,1}$ for each sample $i$triangleright Extract retain probabilities
20:$\mathcal{S}\leftarrow\{i\mid p_{i}>0.5\}$triangleright Default threshold $0.5$
21:return $\{p_{i}\}_{i=1}^{n}$, $\mathcal{S}$
Appendix C Experimental Details
We shift the focus from theoretical analysis and algorithm steps to empirical results. We first detail the training configurations and computational resources used in our experiments.
Mislabeled Sample Detection. For the experiments of simulating mislabeled samples (Sec. 5.1), we fine-tune ViT-B/16 models pretrained on the ImageNet-21k dataset [56]. Training is conducted with a batch size of 32, a learning rate of 1e-4, and the AdamW optimizer. To ensure convergence, we set the training steps to 10 for the DTD and Flowers102 datasets and to 5 for the remaining datasets.
@@ Appendix C Experimental Details
We evaluate Grad-Mimic across various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report performance at $\tau=0.5$ in the main paper (Table 1). Results for other temperature values are provided in Table 9. To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets, using different random seeds for weight initialization. All experiments are conducted on an NVIDIA Tesla A100.
@@
Large-scale Data Curation. For the experiments of multimodal data selection (Sec. 5.2), we follow the training protocol used in DataComp [54, 21] and train CLIP models from scratch using contrastive objective on image-caption pairs. Each model is trained for 5 epochs with a batch size of 4096. The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset.
@@
We analyze different components of the reference model’s weights, specifically the final MLP layers in the image and text encoders. We evaluate Grad-Mimic’s effectiveness using temperature values $\tau$ of 0.03, 0.05, 0.07, 0.3, and 0.5. Performance is assessed across 38 diverse downstream tasks [21]. Results for $\tau=0.5$ and $\tau=0.05$ are reported in the main paper (Table 3), and comprehensive results are provided in Table 8. These experiments are conducted on 8 NVIDIA Tesla A100 GPUs.
Appendix D Comparing to Other Model-based Approaches
We compare Grad-Mimic with other model-based techniques, specifically reference model-based approach and influence function-based approach.
Table 6: Grad-Mimic outperforms other reference model-based methods: Using the last-layer weights of reference model is more effective to prioritize samples for learning while reducing GPU memory usage and substantial computational overhead (see efficiency advantages are in Appendix E).
DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Table 7: Grad-Mimic outperforms other reference model-based methods even when using weaker or distilled reference model as proxy.
CIFAR10 CIFAR100
Noise Level 0.5 0.5
Rho-Loss 93.83 76.05
Using full training set Grad-Mimic 93.92 76.82
Rho-Loss 93.26 75.11
Using hold-out training set Grad-Mimic 93.38 75.81
Figure 6: Grad-Mimic outperforms approximated influence function-based methods: Using reference model weights as selection guide offer advantages on their effectiveness and greater accessibility.
Reference Model-Based Approaches. We compare Grad-Mimic to Rho-Loss [14], a representative method that prioritizes samples based on excess loss, computed as the difference between the loss on the current model $\ell(\theta_{t};s_{i})$ and the loss on a pretrained reference model $\ell(\theta_{\text{ref}};s_{i})$. For a fair comparison, we simulate pretrained reference models using two approaches: (i) training on a noise-free hold-out dataset as used in Rho-Loss [14], and (ii) training on the entire training dataset, consistent with our setup in the main paper. We adopt our mislabeled sample detection setup and evaluate the model performance performance on the testing datasets.
We first demonstrate model performance using reference model that is trained on entire training dataset. As shown in Table 6, Grad-Mimic consistently outperforms Rho-Loss across all datasets and noise levels. Next, we benchmark both methods on CIFAR10 and CIFAR100 with 50% label noise, using a reference model trained on a noise-free hold-out dataset. Such setup simulates more realistic scenarios where the reference model, acting as a distilled or weaker approximation of ideal weights, is used. The results, presented in Table 7, again show that Grad-Mimic achieves higher test accuracy than Rho-Loss.
Moreover, Grad-Mimic offers other computational advantages. Unlike Rho-Loss, which requires loading the entire reference model and performing additional inference passes to compute excess loss, Grad-Mimic scores sample utility directly from discrepancies in model weights—a simpler and more efficient operation. Overall, these results are showing that using only a subset of weights, such as those in the last layer, is sufficient for Grad-Mimic to ensure high performance, further reducing memory usage and improving computational efficiency.
Influence Function-based Approaches. Data influence functions provide a principled way to evaluating the impact of individual training samples on model parameters [57, 58, 59], but their computational cost often limits practical use. Recent studies have explored efficient approximations that estimate sample influence by measuring gradient alignment with a noisy-free validation dataset [19, 20, 18]. We adapt these methods within Grad-Mimic framework by replacing the vector induced by reference model’s weights ($v_{t}$) with gradients computed from a noise-free validation dataset. We evaluate these adaptations on six image datasets by adding 50% label noise and assess the impact of validation set size by varying the number of samples drawn from per class (1, 5, 10, 15, and 20).
Results in Figure 6 show that while larger validation sets enhance data efficiency by enabling models to focus on influential samples, Grad-Mimic, which leverages pretrained reference weights and their induced direction, consistently outperforms influence approximations based on validation datasets. Moreover, relying on labeled validation sets introduces dataset dependencies, incur computational overhead from validation dataset’s gradient computations, and poses challenges in ensuring their quality, as noted in the limitations of GREATS [18]. For example, in the out-of-domain Flowers102 dataset, achieving satisfactory performance required over 2,000 correctly labeled validation samples—yet still fell short of Grad-Mimic’s results.
Appendix E Efficiency Advantages
Figure 7: Grad-Mimic enjoys lower runtime per batch and smaller memory usage than other baselines.
Figure 8: Grad-Mimic leads to faster convergence and improved data utilization.
We provide four distinct lines of evidence for the efficiency gains our technique. They are (i) computational efficiency, (ii) memory usage, (iii) learning efficiency, and (iv) data utilization.
Computational Efficiency and Memory Usage. First, we benchmark per-batch training runtime and peak GPU memory consumption across considered baselines, as presented in Figure 7. Grad-Mimic achieves computational runtime and memory usage comparable to GraNd, relying solely on per-sample training gradients. In contrast, methods like AGRA and approximated influence function-based approaches incur higher computational costs due to additional gradient computations from either another training batch or a noise-free validation dataset. This creates a 2.6 times overhead. Similarly, Grad-Mimic outperforms Rho-Loss in efficiency, as Rho-Loss requires loading a full reference model and performing additional forward passes—substantially increasing compute time. Regarding memory usage, Rho-Loss can exceed all these methods if the reference model size scales significantly.
Learning Efficiency. Second, we evaluate learning efficiency by analyzing convergence curves. As shown in Figure 8 (left), experiments on CIFAR100 with 50% label noise demonstrate that Grad-Mimic converges faster than all compared methods, achieving higher model performance within a fixed number of training epochs. Additionally, Figure 4 (left) illustrates that, on the DataComp medium-scale dataset (over 100 million samples), Grad-Mimic, aided by a reference model, converges faster than standard training, reducing training steps by 20.7%.
Data Utilization. Finally, we introduce a new experiment that measures data efficiency without adding label noise. Instead of simulating mislabeled samples for denoising, we rank clean training samples by their mimic scores and build subsets from the top-kentries. We use a well-trained ViT-B/16 as our reference model and perform Grad-Mimic on clean CIFAR100 set. We evaluate the model performance trained on refined subsets. Results, presented in Figure 8, show that Grad-Mimic’s selected subsets consistently outperform randomly drawn subsets of the same size. Remarkably, with both ViT-B/16 and the larger ViT-L/14, Grad-Mimic matches the full-dataset accuracy while reducing the training data by 20%. This highlights a direct reduction in data-related computational cost.
These findings—lower runtime per batch, reduced memory usage, faster convergence, and improved data utilization—demonstrate that Grad-Mimic achieves significant efficiency gains, fulfilling our motivation.
@@ Appendix E Efficiency Advantages
Appendix F Ablation Studies
@@
Choice of Temperature. We run Grad-Mimic under different temperature values and compare them to baseline methods (Mini-batch SGD and Grad-Match [29]). The results in the simulation experiments are presented in Table 9. In these datasets, we set the noise level to 0.3 and fine-tune ViT-B/16 model under linear probing configuration. Grad-Mimic outperforms both baseline methods at every temperature. Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.
Table 10: Last-layer weights yields the best performance across all datasets.
DTD Oxford-IIIT Pet CIFAR10 CIFAR100
Layer 7 0.575 0.755 0.907 0.752
Layer 8 0.583 0.755 0.908 0.753
Layer 9 0.586 0.740 0.911 0.749
Layer 10 0.582 0.751 0.902 0.753
Layer 11 (last-layer weights) 0.595 0.758 0.912 0.755
Which layer should be mimicked? Next, we study the impact of the layer weights to mimic. We vary the depth of the reference model weights used to obtain the target vector and evaluated the resulting model performance. Results, presented in the Table 10, indicate that mimicking the last-layer weights yields the highest accuracy in most cases. While the optimal choice may depend on architecture or task, these findings suggest that later representations generally offer the richest information and are better choices for estimating sample utility.
Impact of Reference Model Quality. Finally, we investigate how the quality of the reference model impacts the effectiveness of Grad-Mimic, particularly in scenarios where only a weaker or distilled model is available. To simulate such scenarios, we degrade the reference model in two controlled ways:
  • Weight corruption: we inject zero-mean Gaussian noise $(0,\sigma^{2})$ into the reference model’s weights. By varying the variance of the added noise, we analyze the performance of models trained with Grad-Mimic against standard training methods that do not leverage a reference model.
  • Data-limited training: we train the reference model on progressively smaller fractions of its original dataset (shrinking from 1.3B samples to 3.5M samples). Fewer training samples will produce a weaker (or distilled) reference model.
We assess Grad-Mimic in both the mislabeled-sample and multimodal settings, applying each degradation method separately.
We present these two analysis in Figure 9 and Table 11. As expected, increasing noise variance or using less training data degrades the reference model’s performance, which in turn lowers the effectiveness of Grad-Mimic. However, we observe that even under significant noise (heavy degradation), models trained with Grad-Mimic still outperform Mini-batch SGD and can even surpass the performance of the noisy reference model itself (see Figure 9). This highlights that while a high-quality reference model is ideal, a weaker proxy can still provide valuable guidance and improve training over traditional baselines.
Table 11: Grad-Mimic achieves performance gains even when using a weaker reference model.
Scale # of samples to train $\theta_{\text{ref}}$ Training Method ImageNet
ImageNet
dist. shifts
VTAB Retrieval
Average over
38 datasets ($\uparrow$)
Using 1.3B samples Vanilla Training 0.026 0.035 0.139 0.114 0.131
Using 1.3B samples Grad-Mimic 0.026 0.037 0.162 0.118 0.146
Small Using 3.5M samples Grad-Mimic 0.026 0.035 0.158 0.115 0.140
@@
Appendix G Broader Impacts and Limitations
Efficient data selection is a crucial prerequisite for training multimodal models. We demonstrate Grad-Mimic’s effectiveness in identifying and removing low-value samples, which enhances data efficiency and downstream model performance. We do not foresee any direct negative social consequences from the technique itself.
However, if the reference model used to evaluate sample utility is poorly calibrated or biased, there remains a risk of refining low-quality datasets. Appendix F investigates such scenario. It shows that Grad-Mimic remains competitive even when the reference model is heavily degraded, yet some performance loss is unavoidable. We believe that practitioners can easily mitigate this risk by validating the reference model on a small, trusted benchmark before data scoring. This check is easier than the requirements of many recent data-selection methods, which rely on noise-free validation sets. By decoupling data scoring from such hard-to-obtain resources, Grad-Mimic offers greater flexibility and control.